OpenStudy (anonymous):

write as single trigonometric function: sin(90degrees-theta)/cot(90degrees-theta) PLEASE I KNOW THE ANSWER I JUST NEED SOMEONE TO EXPLAIN IT

OpenStudy (misty1212):

\[\sin(x)/\cot(x)=\sin(x)\tan(x)\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (misty1212):

\[\tan(90-x)=-\tan(x)\]

OpenStudy (anonymous):

the even/odd identity

OpenStudy (misty1212):

i am sorry i am wrong \[\sin(90-x)=\cos(x)\]

OpenStudy (misty1212):

and \[\tan(90-x)=\cot(x)\] making this hole thing \[\cos(x)\cot(x)\]